OpenStudy (andrewkaiser333):

In an area in which electricity costs 8 cents/kilowatt-hour, a 5 kW clothes dryer runs for 90 minutes to dry a load of laundry. How much did it cost to dry these clothes? These are the choices $0.20 <------ This is my answer. $0.30 <------ This is my second answer. $0.45 $0.60 Please help!!!!!!!!!!!!

OpenStudy (andrewkaiser333):

@TuringTest @anonymous_user @ganeshie8 @BrianMcGreevy

OpenStudy (andrewkaiser333):

@Schleifspur

OpenStudy (anonymous):

60

OpenStudy (andrewkaiser333):

kk

OpenStudy (anonymous):

calculation of the power your machine needs: P = W * t P = 5 kW * 1.5 h P = 7.5 kWh

OpenStudy (andrewkaiser333):

how did you get it so i know

OpenStudy (anonymous):

5 killowat dryer times 8 cents a killowat for 1.5 hours

OpenStudy (anonymous):

cost calculation: M = P * 8 cents/kWh M = 60 cents

OpenStudy (andrewkaiser333):

ok cool thanks

OpenStudy (andrewkaiser333):

don't know which one to give my medal too i want both of you to have one since yall had it right and helped me

OpenStudy (andrewkaiser333):

i gave one to you schleifspur and if you do not mind since brian helped too you send him a medal please????

OpenStudy (andrewkaiser333):

@Schleifspur